A chemist has 5 gallons of salt solution with a concentration of 0.2 pounds per gallon and another solution with a concentration of 0.5 pounds per gallon. How many gallons of the stronger solution must be added to the weaker solution to get a solution that contains 0.3 pounds per gallon?
Answer by stanbon(75887)

A chemist has 5 gallons of salt solution with a concentration of 0.2 pounds per gallon and another solution with a concentration of 0.5 pounds per gallon. How many gallons of the stronger solution must be added to the weaker solution to get a solution that contains 0.3 pounds per gallon?
------
Equation:
salt + salt = salt
0.2*5 + 0.5x = 0.3(5+x)
---
Multiply thru by 10 to get:
2*5 + 5x = 3*5+3x
2x = 1*5
x = 2.5 gallons (amt of stronger needed for the mixture)
==================
Cheers,
Stan H.